The end of the spring semester means graduation an Number 9
Answer #1

We have:

- x1=-5cm

-magnification=+3 \beta =\frac{x_{2}}{x_{1}}

Thus, x2=-15 cm

Replace in this formula \frac{1}{x_{2}}-\frac{1}{x_{1}}=\frac{1}{f} and get f=+7.5 cm.